### v3

# The low dimensional homology groups of the elementary group of degree two

Abstract.
In this article, we study the first, second, and third homology groups of the elementary group $\mathit{{\rm E}}_{2}(A)$, where $A$ is a commutative ring. In particular, we establish a refined Bloch–Wigner type exact sequence over a semilocal ring (subject to mild restrictions on its residue fields) in which either $1\in({A^{\times}})^{2}$ or $|{A^{\times}}/({A^{\times}})^{2}|\leq 4$.
For a commutative ring $A$ with unit, let $\mathit{{\rm E}}_{2}(A)$ denote the subgroup of the general linear group $\mathit{{\rm GL}}_{2}(A)$ generated by elementary matrices. Clearly, $\mathit{{\rm E}}_{2}(A)\subseteq\mathit{{\rm SL}}_{2}(A)$. We say that $A$ is a ${\rm GE}_{2}(A)$-ring if $\mathit{{\rm E}}_{2}(A)=\mathit{{\rm SL}}_{2}(A)$.
In two previous papers [17] and [18] we studied the low-dimensional homology groups of the elementary group $\mathit{{\rm E}}_{2}(A)$. In this paper, we further investigate these homology groups.
The first main result of this paper concerns the first homology group of $\mathit{{\rm E}}_{2}(A)$. We show that for any commutative ring $A$, there is an exact sequence
$H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow H_{1}(Y_{\bullet}(A^{2}))_{\mathit{{\rm E}}_{2}(A)}\rightarrow A/M\rightarrow H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow 0,$
where $Y_{\bullet}(A^{2})$ is a complex of ${\rm GE}_{2}(A)$-modules obtained from certain unimodular vectors in $A^{2}$ (see Theorem 4.1). Here $M$ is the additive subgroup of $A$ generated by $x(a^{2}-1)$ and $3(b+1)(c+1)$, where $x\in A$ and $a,b,c\in{A^{\times}}$. The rightmost map is induced by the homomorphism $A\rightarrow\mathit{{\rm E}}_{2}(A)$ sending $x$ to the elementary matrix $E_{12}(x)$. The group $A/M$ and its connection with the first homology of $\mathit{{\rm E}}_{2}(A)$ were first studied by Cohn in [3] and [4].
Our second main result concerns the second homology group $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$. For simplicity, let $A$ be a semilocal ring, where all its residue fields have more than three elements. Then it is not difficult to show that this homology group fits into the exact sequence
$H_{2}(\mathrm{B}(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A)\rightarrow 0,$
where $\mathrm{B}(A)$ is the subgroup of $\mathit{{\rm E}}_{2}(A)$ consisting of all matrices whose lower left-hand entry is equal to zero. Here $I(A)$ denotes the fundamental ideal of $A$. More precisely, let $\mathcal{G}_{A}$ be the square class group of $A$ and $\mathcal{I}_{A}$ the augmentation ideal of $\mathbb{Z}[\mathcal{G}_{A}]$. Let $\langle a\rangle$ denotes the element of $\mathcal{G}_{A}$ represented by $a\in{A^{\times}}$. Then we define
$I(A):=\mathcal{I}_{A}/\langle\langle\!\langle a\rangle\!\rangle\langle\!\langle 1-a\rangle\!\rangle:a,1-a\in{A^{\times}}\rangle,$
where $\langle\!\langle a\rangle\!\rangle:=\langle a\rangle-1\in\mathbb{Z}]\mathcal{G}_{A}]$. Observe that $I^{2}(A):=(I(A))^{2}$. Our second result concerns the map $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A)$ appearing in the above exact sequence. When $2\in{A^{\times}}$, we exhibit a cycle whose homology class in $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ maps to the element $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle$, for $a,b\in{A^{\times}}$ (see Proposition 5.2 for the precise statement).
The third main result of this paper is a refined Bloch–Wigner exact sequence over a semilocal domain, which describes the third homology of $\mathit{{\rm E}}_{2}(A)$. More precisely, let $A$ be a semilocal ring such that for any maximal ideal $\mathfrak{m}\in{\rm Specm}(A)$, the residue field $A/\mathfrak{m}$ is either infinite, or if $|A/\mathfrak{m}|=p^{d}$, then (p1)d>6(p-1)d>6. If $1\in{A^{\times}}^{2}$ or $|{A^{\times}}/({A^{\times}})^{2}|\leq 4$, then we obtain the refined Bloch–Wigner exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)}\rightarrow\mathcal{RB}(A)\rightarrow 0,$
where $\mathcal{RB}(A)$ is the refined Bloch group of $A$ (see Theorem 6.3). Here the map $\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)\rightarrow H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ is obtained from the product map
$\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z}),$
which is induced by the map $\mu_{2}(A)\times\mathit{{\rm E}}_{2}(A)\rightarrow\mathit{{\rm E}}_{2}(A)$, $(-1,X)\mapsto-X$. This generalizes the main result of [18]. In [20, Question 0.1], we raised the question of whether, over local rings, the product map
$\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$
is trivial, provided that $|A/\mathfrak{m}_{A}|\neq 2$.
Our main tool for studying the homology groups of $\mathit{{\rm E}}_{2}(A)$ is a first-quadrant spectral sequence derived from the complex $Y_{\bullet}(A^{2})$. All our main results are obtained from a detailed analysis of this spectral sequence.
We now outline the organization of the present paper. In Section 1 we give a brief overview of the groups $\mathit{{\rm E}}_{2}(A)\subseteq\mathit{{\rm SL}}_{2}(A)$, ${\rm GE}_{2}(A)\subseteq\mathit{{\rm GL}}_{2}(A)$, the rank-one Steinberg group $\mathrm{{St}}(2,A)$, and the rank-one $\mathrm{K}_{2}$-group $\mathrm{K}_{2}(2,A)$. In Section 2, we introduce and study an $\mathit{{\rm SL}}_{2}(A)$-complex $L_{\bullet}(A^{2})$ and its $\mathit{{\rm E}}_{2}(A)$-subcomplex $Y_{\bullet}(A^{2})$. In particular, we study the groups $H_{0}$ and $H_{1}$ of these complexes. In Section 3, we introduce and analyze our main spectral sequence. In Section 4, we study the first homology group of $\mathit{{\rm E}}_{2}(A)$ and prove the first exact sequence discussed in this introduction. In Section 5, we study the natural map $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A)$. In Section 6, we establish the refined Bloch–Wigner exact sequence, which describes the third homology of $\mathit{{\rm E}}_{2}(A)$ over semilocal rings.
Notations. Throughout this paper all rings are commutative, except possibly group rings, and possess a unit element $1$. For a commutative ring $A$, let $\mathcal{G}_{A}:={A^{\times}}/({A^{\times}})^{2}$ and define
$\mathcal{W}_{A}:=\{\,a\in A:a(1-a)\in{A^{\times}}\,\}.$
The element of $\mathcal{G}_{A}$ represented by $a\in{A^{\times}}$ is denoted by $\langle a\rangle$, and we set $\langle\!\langle a\rangle\!\rangle:=\langle a\rangle-1\in\mathbb{Z}[\mathcal{G}_{A}]$. For $n\in\mathbb{N}$, let $\mu_{n}(A)$ denote the group of $n$th roots of unity in $A$. Moreover, define
$\mu(A):=\bigcup_{n\in\mathbb{N}}\mu_{n}(A).$
If $\mathcal{B}\rightarrow\mathcal{A}$ is a homomorphism of abelian groups, we write $\mathcal{A}/\mathcal{B}$ for ${\rm coker}(\mathcal{B}\rightarrow\mathcal{A})$.
1. Elementary and special linear groups of degree two
Let $A$ be a commutative ring. Let $\mathit{{\rm E}}_{2}(A)$ be the subgroup of $\mathit{{\rm GL}}_{2}(A)$ generated by the elementary matrices $E_{12}(a):=\begin{pmatrix}1&a\\ 0&1\end{pmatrix}$ and $E_{21}(a):=\begin{pmatrix}1&0\\ a&1\end{pmatrix}$, $a\in A$. Clearly
$\mathit{{\rm E}}_{2}(A)\subseteq\mathit{{\rm SL}}_{2}(A).$
For most commutative rings, $\mathit{{\rm E}}_{2}(A)$ is a proper subgroup of $\mathit{{\rm SL}}_{2}(A)$ and most of times it is even not a normal subgroup. We say that $A$ is a ${\rm GE}_{2}$-ring if
$\mathit{{\rm E}}_{2}(A)=\mathit{{\rm SL}}_{2}(A).$
Let ${\rm GE}_{2}(A)$ be the subgroup of $\mathit{{\rm GL}}_{2}(A)$ generated by diagonal and elementary matrices. It is easy to verify that $\mathit{{\rm E}}_{2}(A)$ is normal in ${\rm GE}_{2}(A)$ and the center of ${\rm GE}_{2}(A)$ is ${A^{\times}}I_{2}$. Note that $A$ is a ${\rm GE}_{2}$-ring if and only if
${\rm GE}_{2}(A)=\mathit{{\rm GL}}_{2}(A).$
Semilocal rings and euclidean domains are examples of ${\rm GE}_{2}$-rings [23, p. 245], [3, §2]. If $F$ is a field, then $F[X_{1},\dots,X_{d}]$, $d>1$, is not a ${\rm GE}_{2}$-ring [3, page 26].
A quadratic ring is a subring of $\mathbb{C}$ of the form $\mathbb{Z}[\omega]:=\mathbb{Z}\oplus\mathbb{Z}\omega$, where $\omega$ is the root of an irreducible quadratic polynomial of the form $X^{2}\pm d$ or $X^{2}+X\pm d$ with $d$ a positive integer. The quadratic rings are split by the choice of sign into real quadratic rings $\mathbb{Z}[\sqrt{d}]$, $\mathbb{Z}[\frac{1}{2}(1+\sqrt{1+4d})]$ and imaginary quadratic rings $\mathbb{Z}[\sqrt{-d}]$, $\mathbb{Z}[\frac{1}{2}(1+\sqrt{1-4d})]$.
Theorem 1.1.
Let $A=\mathbb{Z}[\sqrt{-d}]$ or $A=\mathbb{Z}[\frac{1}{2}(1+\sqrt{1-4d})]$, where $d$ is a positive integer.
(i) If $d\geq 4$, then $\mathit{{\rm E}}_{2}(A)$ is a non-normal, infinite-index subgroup of $\mathit{{\rm SL}}_{2}(A)$.
(ii) If $1\leq d<4$, then $\mathit{{\rm E}}_{2}(A)=\mathit{{\rm SL}}_{2}(A)$, i.e. $A$ is a ${\rm GE}_{2}$-ring.
Proof.
See [22, Theorem 1.5]. ∎
If $1\leq d<4$, then $A$ is one of the rings $\mathbb{Z}[\sqrt{-1}]$, $\mathbb{Z}[\sqrt{-2}]$, $\mathbb{Z}[\sqrt{-3}]$, $\mathbb{Z}[\frac{1}{2}(1+\sqrt{-3})]$, $\mathbb{Z}[\frac{1}{2}(1+\sqrt{-7})]$, $\mathbb{Z}[\frac{1}{2}(1+\sqrt{-11})]$. Thus these rings are ${\rm GE}_{2}$-rings.
Let $K$ be an algebraic number field and let $\mathcal{O}_{K}$ be its ring of algebraic integers. An order of $K$ is a subring $\mathcal{O}$ of $\mathcal{O}_{K}$ which is a free $\mathbb{Z}$-module of rank $n=[K:\mathbb{Q}]$. Thus any quadratic ring is an order.
Theorem 1.2 (Liehl [14]).
Let A be an order in a number field which is not an imaginary quadratic ring. Then $\mathit{{\rm E}}_{2}(A)$ is a normal, finite-index subgroup in $\mathit{{\rm SL}}_{2}(A)$. Moreover, if the number field has a real embedding, then $\mathit{{\rm E}}_{2}(A)=\mathit{{\rm SL}}_{2}(A)$.
Proof.
See [22, Theorem 1.6]. ∎
The rank one Steinberg group $\mathrm{{St}}(2,A)$ of a commutative ring $A$ is the group with generators $x_{12}(r)$ and $x_{21}(s)$, $r,s\in A$, subject to the Steinberg relations
  • ($\alpha$)
    $x_{ij}(r)x_{ij}(s)=x_{ij}(r+s)$ for any $r,s\in A$,
  • ($\beta$)
    $w_{ij}(u)x_{ji}(r)w_{ij}(u)^{-1}=x_{ij}(-u^{2}r)$, for any $u\in{A^{\times}}$ and $r\in A$,
    where $w_{ij}(u):=x_{ij}(u)x_{ji}(-u^{-1})x_{ij}(u)$.
The natural map $\mathrm{{St}}(2,A)\rightarrow\mathit{{\rm E}}_{2}(A)$, given by $x_{ij}(r)\mapsto E_{ij}(r)$, is a well defined homomorphism of groups. The kernel of this map is called the rank one $K_{2}$-group of $A$ and is denoted by $\mathrm{K}_{2}(2,A)$:
$\mathrm{K}_{2}(2,A):=\ker(\mathrm{{St}}(2,A)\rightarrow\mathit{{\rm E}}_{2}(A)).$
For $u\in{A^{\times}}$, let
$h_{ij}(u):=w_{ij}(u)w_{ij}(-1).$
An element of the form
$\{v,u\}:=\{v,u\}_{12}=h_{12}(uv)h_{12}^{-1}(u)h_{12}(v)^{-1}$
is called a Steinberg symbol. This is an element of $\mathrm{K}_{2}(2,A)$ and is a central element of $\mathrm{{St}}(2,A)$ (see [6, §9]).
Let $\mathrm{C}(2,A)$ be the subgroup of $\mathrm{K}_{2}(2,A)$ generated by the Steinberg symbols. A ring $A$ is called universal for ${\rm GE}_{2}$ if $\mathrm{K}_{2}(2,A)$ is generated by Steinberg symbols [6, p. 251]. This definition coincides with the original definition of a ring universal for ${\rm GE}_{2}$ given by Cohn in [3, page 8] (for a proof of this fact see [12, App. A]).
2. The complex of unimodular vectors
A vector ${\boldsymbol{u}}={\left(\!\!\!\begin{array}[]{cc}u_{1}\\ u_{2}\end{array}\!\!\!\right)}\in A^{2}$ is called unimodular if there exists a vector ${\boldsymbol{v}}={\left(\!\!\!\begin{array}[]{cc}v_{1}\\ v_{2}\end{array}\!\!\!\right)}\in A^{2}$ such that $({\boldsymbol{u}},{\boldsymbol{v}}):={\left(\!\begin{array}[]{cc}\!\!u_{1}&\!\!v_{1}\\ \!\!u_{2}&\!\!v_{2}\end{array}\!\!\!\right)}\in\mathit{{\rm GL}}_{2}(A)$ and it is called ${\rm GE}_{2}$-unimodular if $({\boldsymbol{u}},{\boldsymbol{v}})\in{\rm GE}_{2}(A)$.
For any non-negative integer $n$, let $L_{n}(A^{2})$ be the free abelian group generated by the set of all $(n+1)$-tuples $(\langle{\boldsymbol{v}_{0}}\rangle,\dots,\langle{\boldsymbol{v}_{n}}\rangle)$, where every vector ${\boldsymbol{v}_{i}}\in A^{2}$ is unimodular and for any $i\neq j$, $({\boldsymbol{v}_{i}},{\boldsymbol{v}_{j}})\in\mathit{{\rm GL}}_{2}(A)$. Let $Y_{n}(A^{2})$ be the free abelian subgroup of $L_{n}(A^{2})$ generated by the set of all $(n+1)$-tuples $(\langle{\boldsymbol{v}_{0}}\rangle,\dots,\langle{\boldsymbol{v}_{n}}\rangle)$, where every vector ${\boldsymbol{v}_{i}}\in A^{2}$ is ${\rm GE}_{2}$-unimodular. It is easy to see if $\boldsymbol{u}$} is ${\rm GE}_{2}$-unimodular and $({\boldsymbol{u}},{\boldsymbol{v}})\in\mathit{{\rm GL}}_{2}(A)$, then $({\boldsymbol{u}},{\boldsymbol{v}})\in{\rm GE}_{2}(A)$ and hence $\boldsymbol{v}$} is ${\rm GE}_{2}$-unimodular.
We consider $L_{n}(A^{2})$ (resp. $Y_{n}(A^{2})$) as a left $\mathit{{\rm SL}}_{2}(A)$-module (resp. left $\mathit{{\rm E}}_{2}(A)$-module) in a natural way. If necessary, we convert this action to a right action by the definition $m.g:=g^{-1}m$. Let us define the $n$-th differential operator
$\partial_{n}^{L}:L_{n}(A^{2})\rightarrow L_{n-1}(A^{2}),\ \ n\geq 1,$
as an alternating sum of face operators which throws away the $i$-th component of generators. Hence we have the complex $L_{\bullet}(A^{2})$ with the $\mathit{{\rm E}}_{2}(A)$-subcomplex $Y_{\bullet}(A^{2})$.
For a subgroup $H$ of a group $G$ and any $H$-module $M$, let ${\rm Ind}_{H}^{G}M:=\mathbb{Z}[G]\otimes_{H}M$ be the induced module (from $H$ to $G$). This extension of scalars is called induction from $H$ to $G$ [2, §5, Chap. III].
Proposition 2.1.
For any commutative ring $A$, $L_{\bullet}(A^{2})\simeq{\rm Ind}_{\mathit{{\rm E}}_{2}(A)}^{\mathit{{\rm SL}}_{2}(A)}Y_{\bullet}(A^{2})$.
Proof.
Clearly the map
$\varphi_{\bullet}:\mathbb{Z}[\mathit{{\rm SL}}_{2}(A)]\otimes_{\mathit{{\rm E}}_{2}(A)}Y_{\bullet}(A^{2})\rightarrow L_{\bullet}(A^{2})$
defined by
$g\otimes(\langle\boldsymbol{v_{0}}\rangle,\langle\boldsymbol{v_{1}}\rangle,\dots,\langle\boldsymbol{v_{n}}\rangle)\mapsto(\langle g\boldsymbol{v_{0}}\rangle,\langle g\boldsymbol{v_{1}}\rangle,\dots,\langle g\boldsymbol{v_{n}}\rangle),$
is a well-defined morphism of $\mathit{{\rm SL}}_{2}(A)$-complexes. This is an isomorphism with the inverse morphism
$\psi_{\bullet}:L_{\bullet}(A^{2})\rightarrow\mathbb{Z}[\mathit{{\rm SL}}_{2}(A)]\otimes_{\mathit{{\rm E}}_{2}(A)}Y_{\bullet}(A^{2})$
defined by
$(\langle\boldsymbol{v_{0}}\rangle,\langle\boldsymbol{v_{1}}\rangle,\dots,\langle\boldsymbol{v_{n}}\rangle)\mapsto g\otimes(\langle\boldsymbol{e_{1}}\rangle,\langle g^{-1}\boldsymbol{v_{1}}\rangle,\dots,\langle g^{-1}\boldsymbol{v_{n}}\rangle),$
where $g\boldsymbol{e_{1}}=\boldsymbol{v_{0}}$ for some $g\in\mathit{{\rm SL}}_{2}(A)$. ∎
Corollary 2.2.
For any non-negative integer $n$, $H_{n}(L_{\bullet}(A^{2}))\simeq{\rm Ind}_{\mathit{{\rm E}}_{2}(A)}^{\mathit{{\rm SL}}_{2}(A)}H_{n}(Y_{\bullet}(A^{2}))$.
Proof.
Since the functor ${\rm Ind}_{\mathit{{\rm E}}_{2}(A)}^{\mathit{{\rm SL}}_{2}(A)}$ is exact on the category of $\mathbb{Z}[\mathit{{\rm E}}_{2}(A)]$-modules (since $\mathbb{Z}[\mathit{{\rm SL}}_{2}(A)]$ is a free $\mathbb{Z}[\mathit{{\rm E}}_{2}(A)]$-module), the claim follows from the previous lemma. ∎
Let $\partial_{-1}=\varepsilon:L_{0}(A^{2})\rightarrow\mathbb{Z}$ be the natural augmentation map, which is given by $\sum_{i}n_{i}(\langle{\boldsymbol{v}_{0,i}}\rangle)\mapsto\sum_{i}n_{i}$. Then we have the complexes $L_{\bullet}(A^{2})\overset{\varepsilon}{\rightarrow}\mathbb{Z}$ and $Y_{\bullet}(A^{2})\overset{\varepsilon}{\rightarrow}\mathbb{Z}$.
Theorem 2.3 (Hutchinson).
For any commutative ring $A$, $H_{0}(Y_{\bullet}(A^{2}))\overset{\bar{\varepsilon}}{\simeq}\mathbb{Z}$, where $\bar{\varepsilon}$ is induced by $\varepsilon$. In other words, the complex $Y_{1}(A^{2})\overset{\partial_{1}^{Y}}{\longrightarrow}Y_{0}(A^{2})\overset{\varepsilon}{\rightarrow}\mathbb{Z}\rightarrow 0$ always is exact. Moreover,
$H_{0}(L_{\bullet}(A^{2}))\simeq{\rm Ind}_{\mathit{{\rm E}}_{2}(A)}^{\mathit{{\rm SL}}_{2}(A)}\mathbb{Z}\simeq\mathbb{Z}[\mathit{{\rm SL}}_{2}(A)/\mathit{{\rm E}}_{2}(A)],$
where the right side is the coset module.
Proof.
For the first isomorphism see the proof of [12, Theorem 3.3]. The isomorphism $H_{0}(L_{\bullet}(A^{2}))\simeq{\rm Ind}_{\mathit{{\rm E}}_{2}(A)}^{\mathit{{\rm SL}}_{2}(A)}\mathbb{Z}$ follows from the first isomorphism and Corollary 2.2. The other isomorphism follows from the fact that $\mathbb{Z}[\mathit{{\rm SL}}_{2}(A)]$ is a free $\mathbb{Z}[\mathit{{\rm E}}_{2}(A)]$-module with a basis consisting of one representative for each element of the set $\mathit{{\rm SL}}_{2}(A)/\mathit{{\rm E}}_{2}(A)$.
Example 2.4.
(i) Let $A=\mathbb{Z}[\sqrt{-d}]$ or $A=\mathbb{Z}[\frac{1}{2}(1+\sqrt{1-4d})]$, where $d\geq 4$. Then by Theorem 2.3 and Theorem 1.1, $H_{0}(L_{\bullet}(A^{2}))$ is a free abelian group of infinite and countable rank.
(ii) If $A$ is an order in a number field which is not an imaginary quadratic ring, then $\mathit{{\rm E}}_{2}(A)$ is a normal and finite-index subgroup of $\mathit{{\rm SL}}_{2}(A)$ (see Theorem 1.2). Hence $H_{0}(L_{\bullet}(A^{2}))$ is a free abelian group of finite rank.
Theorem 2.5 (Hutchinson).
For any commutative ring $A$, we have the isomorphisms
$H_{1}(Y_{\bullet}(A^{2}))\simeq\bigg(\displaystyle{\frac{\mathrm{K}_{2}(2,A)}{\mathrm{C}(2,A)}}\bigg)^{\rm ab},\ \ \ \ \ H_{1}(L_{\bullet}(A^{2}))\simeq{\rm Ind}_{\mathit{{\rm E}}_{2}(A)}^{\mathit{{\rm SL}}_{2}(A)}\bigg(\displaystyle{\frac{\mathrm{K}_{2}(2,A)}{\mathrm{C}(2,A)}}\bigg)^{\rm ab}.$
In particular if $A$ is universal for ${\rm GE}_{2}$, then $Y_{\bullet}(A^{2})\overset{\epsilon}{\rightarrow}\mathbb{Z}$ and $L_{\bullet}(A^{2})\overset{\epsilon}{\rightarrow}\mathbb{Z}$ are exact in dimension $1$.
Proof.
The first isomorphism follows from [12, Theorem 7.2]. The second isomorphism follows from the first isomorphism and Proposition $2.$ ∎
It follows from Lemma 2.1 and Shapiro’s Lemma that the inclusion $Y_{\bullet}(A^{2})\rightarrow L_{\bullet}(A^{2})$ induces isomorphisms of the homology groups
$H_{q}(\mathit{{\rm E}}_{2}(A),Y_{p}(A^{2}))\simeq H_{q}(\mathit{{\rm SL}}_{2}(A),L_{p}(A^{2}))$
for all $p,q$. For any $n\geq 0$, let
$Z_{n}^{Y}(A^{2}):=\ker(\partial_{n}^{Y}),\ \ \ \ \ Z_{n}^{L}(A^{2}):=\ker(\partial_{n}^{L}).$
Lemma 2.6.
For all $p,q\geq 0$, we have the isomorphism
$H_{q}(\mathit{{\rm E}}_{2}(A),Z_{p}^{Y}(A^{2}))\simeq H_{q}(\mathit{{\rm SL}}_{2}(A),Z_{p}^{L}(A^{2})).$
Proof.
The functor ${\rm Ind}_{\mathit{{\rm E}}_{2}(A)}^{\mathit{{\rm SL}}_{2}(A)}$ is exact on the category of $\mathbb{Z}[\mathit{{\rm E}}_{2}(A)]$-modules. Thus we have the isomorphism
$Z_{n}^{L}(A^{2})\simeq{\rm Ind}_{\mathit{{\rm E}}_{2}(A)}^{\mathit{{\rm SL}}_{2}(A)}Z_{n}^{Y}(A^{2}).$
Now the claim follows by applying Shapiro’s Lemma to this isomorphism. ∎
Let $G$ be a group and let $C_{\bullet}$ be a complex of left $G$-modules: $C_{\bullet}:\ \ \cdots\rightarrow C_{1}\rightarrow C_{0}\rightarrow 0$. The $n$-th homology of $G$ with coefficients in $C_{\bullet}$, denoted by
$H_{n}(G,C_{\bullet}),$
is defined as the $n$-th homology of the total complex of the double complex $F_{\bullet}\otimes_{G}C_{\bullet}$, where $F_{\bullet}\rightarrow\mathbb{Z}$ is a projective resolution of $\mathbb{Z}$ over $G$ [2, §3, Chap. VII].
By applying Shapiro’s Lemma to the isomorphism of Proposition 2.1, for any non-negative integer $n$, we obtain the isomorphism
$H_{n}(\mathit{{\rm E}}_{2}(A),Y_{\bullet}(A^{2}))\simeq H_{n}(\mathit{{\rm SL}}_{2}(A),L_{\bullet}(A^{2})).$
Let $\mathbb{Z}_{\bullet}$ be the complex with terms $\mathbb{Z}_{0}=\mathbb{Z}$ and $\mathbb{Z}_{n}=0$ for $n\neq 0$. Then $\varepsilon$ induces the natural commutative diagram of complexes
$Y_{\bullet}(A^{2})$$L_{\bullet}(A^{2})$$\mathbb{Z}_{\bullet}$$\scriptstyle{\varepsilon}$$\scriptstyle{\varepsilon}$
From this diagram we obtain the commutative diagram
[…]
In Theorem 4.1 below we give a satisfactory description of $H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ over any commutative ring, while the study of $H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$ seems much harder (see for example [24] and [7]).
3. The main spectral sequence and the homology of $\mathit{{\rm E}}_{2}$
Let $G$ be a group and let $C_{\bullet}$ be a complex of left $G$-modules. Let $B_{\bullet}(G)\rightarrow\mathbb{Z}$ be the bar resolution of $G$. From the double complex
$B_{\bullet}(G)\otimes_{G}C_{\bullet}$
we obtain two spectral sequences
$\mathsf{E}_{p,q}^{2}(G)=H_{p}(G,H_{q}(C_{\bullet}))\Rightarrow H_{p+q}(G,C_{\bullet}),$
and
$\mathrm{E}_{p,q}^{1}(G)=H_{q}(G,C_{p})\Rightarrow H_{p+q}(G,C_{\bullet}),$
(see [2, §5, Chap. VII]).
Lemma 3.1.
Let the complex $C_{\bullet}$ be exact for $1\leq i\leq n$ and $M:=H_{0}(C_{\bullet})$. Then $H_{i}(G,C_{\bullet})\simeq H_{i}(G,M)$ for $0\leq i\leq n$.
Proof.
Apply [2, Chap. VII, Proposition 5.2] to the morphism of complexes $C_{\bullet}\rightarrow M_{\bullet}$, where $M_{\bullet}$ is the complex consisting of the single module $M$ concentrated in degree 0. ∎
Thus if $C_{\bullet}$ is exact for every $i\geq 1$, then we have the spectral sequence
(3.1) $\mathrm{E}_{p,q}^{1}(G)=H_{q}(G,C_{p})\Rightarrow H_{p+q}(G,H_{0}(C_{\bullet})).$
For any commutative ring $A$, the sequence
$0\rightarrow Z_{1}^{Y}(A^{2})\overset{{\rm inc}}{\rightarrow}Y_{1}(A^{2})\rightarrow Y_{0}(A^{2})\overset{\varepsilon}{\rightarrow}\mathbb{Z}\rightarrow 0,$
is exact (Theorem 2.3). Thus by (3.1) we have the first quadrant spectral sequence
$E^{1}_{p.q}=\left\{\begin{array}[]{ll}H_{q}(\mathit{{\rm E}}_{2}(A),Y_{p}(A^{2}))&p=0,1\\ H_{q}(\mathit{{\rm E}}_{2}(A),Z_{1}^{Y}(A^{2}))&p=2\\ 0&p>2\end{array}\right.\Longrightarrow H_{p+q}(\mathit{{\rm E}}_{2}(A),\mathbb{Z}).$
The group $\mathit{{\rm E}}_{2}(A)$ acts transitively on the sets of generators of $Y_{i}(A^{2})$ for $i=0,1$. Let
${\boldsymbol{\infty}}:=\langle{\boldsymbol{e}_{1}}\rangle,\ \ \ {\boldsymbol{0}}:=\langle{\boldsymbol{e}_{2}}\rangle,\ \ \ {\boldsymbol{a}}:=\langle{\boldsymbol{e}_{1}}+a{\boldsymbol{e}_{2}}\rangle,\ \ \ a\in{A^{\times}},$
[…]
$Y_{0}(A^{2})\simeq{\rm Ind}_{\mathrm{B}(A)}^{\mathit{{\rm E}}_{2}(A)}\mathbb{Z},\ \ \ \ \ \ \ \ \ \ \ \ Y_{1}(A^{2})\simeq{\rm Ind}_{\mathrm{T}(A)}^{\mathit{{\rm E}}_{2}(A)}\mathbb{Z},$
where
$\mathrm{B}(A):={\rm Stab}_{\mathit{{\rm E}}_{2}(A)}({\boldsymbol{\infty}})=\Bigg\{\begin{pmatrix}a&b\\ 0&a^{-1}\end{pmatrix}:a\in{A^{\times}},b\in A\bigg\},$
$\mathrm{T}(A):={\rm Stab}_{\mathit{{\rm E}}_{2}(A)}({\boldsymbol{\infty}},{\boldsymbol{0}})=\Big\{D(a):a\in{A^{\times}}\Big\},$
where $D(a):=\begin{pmatrix}a&0\\ 0&a^{-1}\end{pmatrix}$ (see [2, §3, Chap. I]). Note that $\mathrm{T}(A)\simeq{A^{\times}}$ and $\mathrm{B}(A)\simeq A\rtimes{A^{\times}}$. In our calculations sometimes we identify $\mathrm{T}(A)$ with $A^{\times}$}. By Shapiro’s lemma we have
$\mathrm{E}_{0,q}^{1}\simeq H_{q}(\mathrm{B}(A),\mathbb{Z}),\ \ \ \ \ \ \mathrm{E}_{1,q}^{1}\simeq H_{q}(\mathrm{T}(A),\mathbb{Z}).$
In particular, $\mathrm{E}_{0,0}^{1}\simeq\mathbb{Z}\simeq\mathrm{E}_{1,0}^{1}$. Moreover, $d_{1,q}^{1}=H_{q}(\sigma)-H_{q}({\rm inc})$, where $\sigma:\mathrm{T}(A)\rightarrow\mathrm{B}(A)$ is given by $\sigma(D(a))=wD(a)w^{-1}=D(a^{-1})$ for $w={\left(\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!0\end{array}\!\!\!\right)}$. This easily implies that $d_{1,0}^{1}$ is trivial and
$d_{1,1}^{1}:\mathrm{T}(A)\rightarrow\mathrm{B}(A)^{\rm ab}=H_{1}(\mathrm{B}(A),\mathbb{Z}),\ \ \ \ \ D(a)\mapsto D(a)^{-2}.$
In fact, note that
$\mathrm{B}(A)^{\rm ab}\simeq\mathrm{T}(A)\oplus A_{A^{\times}}\simeq{A^{\times}}\oplus A_{A^{\times}},$
where
$A_{A^{\times}}=A/\langle a^{2}-1:a\in{A^{\times}}\rangle.$
Thus $d_{1,1}^{1}(D(a))=(D(a)^{-2},0)$. This implies that
$\mathrm{E}_{0,1}^{2}\simeq\mathcal{G}_{A}\oplus A_{A^{\times}},\ \ \ \ker(d_{1,1}^{1})\simeq\mu_{2}(A).$
Furthermore, it is not difficult to see that $d_{2,1}^{1}:H_{1}(\mathit{{\rm E}}_{2}(A),Z_{1}^{Y}(A^{2}))\rightarrow\mu_{2}(A)$ is surjective. In fact, for any $b\in\mu_{2}(A)$, we have $d_{2,1}^{1}([b]\otimes\partial_{2}^{Y}({\boldsymbol{\infty}},{\boldsymbol{0}},{\boldsymbol{a}}))=b$. Therefore
$\mathrm{E}_{1,1}^{2}=0.$
The orbits of the action of $\mathit{{\rm E}}_{2}(A)$ on $Y_{2}(A)$ are represented by
$\langle a\rangle[\ ]:=({\boldsymbol{\infty}},{\boldsymbol{0}},{\boldsymbol{a}}),\ \ \ \ \langle a\rangle\in\mathcal{G}_{A}.$
Therefore $Y_{2}(A^{2})_{\mathit{{\rm E}}_{2}(A)}\simeq\mathbb{Z}[\mathcal{G}_{A}]$. Moreover, the map
$\mathbb{Z}[\mathcal{G}_{A}]\simeq Y_{2}(A^{2})_{\mathit{{\rm E}}_{2}(A)}\rightarrow Y_{1}(A^{2})_{\mathit{{\rm E}}_{2}(A)}\simeq\mathbb{Z}$
coincides with the augmentation map. Now from the composite
$\mathbb{Z}[\mathcal{G}_{A}]\simeq Y_{2}(A^{2})_{\mathit{{\rm E}}_{2}(A)}\rightarrow Z_{1}^{Y}(A^{2})_{\mathit{{\rm E}}_{2}(A)}\rightarrow Y_{1}(A^{2})_{\mathit{{\rm E}}_{2}(A)}\simeq\mathbb{Z}$
it follows that $d_{2,0}^{1}:H_{0}(\mathit{{\rm E}}_{2}(A),Z_{1}^{Y}(A^{2}))\rightarrow\mathbb{Z}$ is surjective. Hence $\mathrm{E}_{1,0}^{2}=0$. Let
${\rm GW}(A):=H_{0}(\mathit{{\rm E}}_{2}(A),Z_{1}^{Y}(A^{2})),\ \ \ \ \epsilon:=d_{2,0}^{1}:{\rm GW}(A)\rightarrow\mathbb{Z}.$
We denote the kernel of $\epsilon$ by $I(A)$.
Combining all these results, we obtain that the first page of the spectral sequence has the following form:
[…]
The second page of the spectral sequence is given by:
[…]
Let
$\overline{{\rm GW}}(A):=\mathbb{Z}[\mathcal{G}_{A}]/\langle\langle\!\langle a\rangle\!\rangle\langle\!\langle 1-a\rangle\!\rangle:a\in\mathcal{W}_{A}\rangle.$
The augmentation map $\mathbb{Z}[\mathcal{G}_{A}]\rightarrow\mathbb{Z}$ induces the natural map $\bar{\epsilon}:\overline{{\rm GW}}(A)\rightarrow\mathbb{Z}$. The kernel of this map is denoted by $\bar{I}(A)$. Note that
$\bar{I}(A):=\mathcal{I}_{A}/\langle\langle\!\langle a\rangle\!\rangle\langle\!\langle 1-a\rangle\!\rangle:a\in\mathcal{W}_{A}\rangle,$
where $\mathcal{I}_{A}$ is the augmentation ideal of $\mathbb{Z}[\mathcal{G}_{A}]$.
From the exact sequence $0\rightarrow B_{1}^{Y}(A^{2})\rightarrow Z_{1}^{Y}(A^{2})\rightarrow H_{1}(Y_{\bullet}(A^{2}))\rightarrow 0$, we obtain the exact sequence
$H_{0}(\mathit{{\rm E}}_{2}(A),B_{1}^{Y}(A^{2}))\rightarrow{\rm GW}(A)\rightarrow H_{1}(Y_{\bullet}(A^{2}))_{\mathit{{\rm E}}_{2}(A)}\rightarrow 0.$
From the sequence $Y_{3}(A^{2})\rightarrow Y_{2}(A^{2})\twoheadrightarrow B_{1}^{Y}(A^{2})$ we obtain the sequence
$H_{0}(\mathit{{\rm E}}_{2}(A),Y_{3}(A^{2}))\rightarrow H_{0}(\mathit{{\rm E}}_{2}(A),Y_{2}(A^{2}))\twoheadrightarrow H_{0}(\mathit{{\rm E}}_{2}(A),B_{1}^{Y}(A^{2})).$
We showed in above that $Y_{2}(A^{2})_{\mathit{{\rm E}}_{2}(A)}\simeq\mathbb{Z}[\mathcal{G}_{A}]$. The orbits of the action of $\mathit{{\rm E}}_{2}(A)$ on $Y_{3}(A)$ are represented by
$\langle a\rangle[x]:=({\boldsymbol{\infty}},{\boldsymbol{0}},{\boldsymbol{a}},\boldsymbol{ax}),\ \ \langle a\rangle\in\mathcal{G}_{A},x\in\mathcal{W}_{A}.$
Thus $Y_{3}(A^{2})_{\mathit{{\rm E}}_{2}(A)}$ is the free $\mathbb{Z}[\mathcal{G}_{A}]$-module generated by the symbols $[x]$, $x\in\mathcal{W}_{A}$. Moreover, the map
$Y_{3}(A^{2})_{\mathit{{\rm E}}_{2}(A)}\rightarrow Y_{2}(A^{2})_{\mathit{{\rm E}}_{2}(A)}\simeq\mathbb{Z}[\mathcal{G}_{A}]$
is given by $x]\mapsto\langle\!\langle x\rangle\!\rangle\langle\!\langle 1-x\rangle\!\rangle$. All these induce a surjective map
$\overline{{\rm GW}}(A)\twoheadrightarrow H_{0}(\mathit{{\rm E}}_{2}(A),B_{1}^{Y}(A^{2})).$
Thus we have the exact sequence
$\overline{{\rm GW}}(A)\rightarrow{\rm GW}(A)\rightarrow H_{1}(Y_{\bullet}(A^{2}))_{\mathit{{\rm E}}_{2}(A)}\rightarrow 0.$
Finally from this we obtain the exact sequence
(3.2) $\bar{I}(A)\rightarrow I(A)\rightarrow H_{1}(Y_{\bullet}(A^{2}))_{\mathit{{\rm E}}_{2}(A)}\rightarrow 0.$
4. The first homology of $\mathit{{\rm E}}_{2}$
In this section, we show that the first homology group of $\mathit{{\rm E}}_{2}(A)$ fits into a natural exact sequence that almost completely describes its structure.
Theorem 4.1.
Let $A$ be a commutative ring and $M$ the additive subgroup of $A$ generated by $x(a^{2}-1)$ and $3(b+1)(c+1)$, where $x\in A$ and $a,b,c\in{A^{\times}}$. Then there is an exact sequence of $\mathcal{G}_{A}$-modules
$H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow H_{1}(Y_{\bullet}(A^{2}))_{\mathit{{\rm E}}_{2}(A)}\rightarrow A/M\rightarrow H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow 0.$
Proof.
By an easy analysis of the above second page of the spectral sequence $E_{\bullet,\bullet}$, we obtain the exact sequence
$H_{2}(\mathrm{B}(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow I(A)\overset{d_{2,0}^{2}}{\longrightarrow}\mathcal{G}_{A}\oplus A_{A^{\times}}\rightarrow H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow 0.$
Let’s study the composite map $\bar{I}(A)\rightarrow I(A)\overset{d_{2,0}^{2}}{\longrightarrow}\mathcal{G}_{A}\oplus A_{A^{\times}}$. Take the element $\langle\!\langle a\rangle\!\rangle\in\bar{I}(A)$, $a\in{A^{\times}}$. This is represented by
$X_{a}:=[\ ]\otimes\Big((\boldsymbol{\infty},\boldsymbol{0},\boldsymbol{a})-(\boldsymbol{\infty},\boldsymbol{0},\boldsymbol{1})\Big)\in B_{0}(\mathit{{\rm E}}_{2}(A))\otimes_{\mathit{{\rm E}}_{2}(A)}Y_{2}(A^{2}).$
By passing to $I(A)$, the image of $\langle\!\langle a\rangle\!\rangle$ is represented by
[…]
which we denote again by $X_{a}$. Thus $d^{2}_{2,0}(\langle\!\langle a\rangle\!\rangle)=d^{2}_{2,0}(\overline{X_{a}})$. To calculate the above composition we look at the diagram
$B_{1}(\mathit{{\rm E}}_{2}(A))\otimes Y_{0}(A^{2})$$B_{1}(\mathit{{\rm E}}_{2}(A))\otimes Y_{1}(A^{2})$$B_{0}(\mathit{{\rm E}}_{2}(A))\otimes Y_{1}(A^{2})$$B_{0}(\mathit{{\rm E}}_{2}(A))\otimes Z_{1}(A^{2}).$$\scriptstyle{{\rm id}\otimes\partial_{1}}$$\scriptstyle{d_{1}\otimes{\rm id}}$$\scriptstyle{{\rm id}\otimes{\rm inc}}$
We have
$({\rm id}_{B_{0}(\mathit{{\rm E}}_{2}(A))}\otimes{\rm inc})(X_{a})$ $=[\ ]\otimes\Big((\boldsymbol{0},\boldsymbol{a})-(\boldsymbol{\infty},\boldsymbol{a})-(\boldsymbol{0},\boldsymbol{1})+(\boldsymbol{\infty},\boldsymbol{1})\Big)$
$=[\ ]\otimes\Big(g_{a}(\boldsymbol{\infty},\boldsymbol{0})-h_{a}(\boldsymbol{\infty},\boldsymbol{0})-g_{1}(\boldsymbol{\infty},\boldsymbol{0})+h_{1}(\boldsymbol{\infty},\boldsymbol{0})\Big)$
$=(g^{-1}_{a}-h^{-1}_{a}-g^{-1}_{1}+h^{-1}_{1})[\ ]\otimes(\boldsymbol{\infty},\boldsymbol{0}),$
where $g_{a}=\left(\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!a\end{array}\!\!\!\right)$, $h_{a}=\left(\!\begin{array}[]{cc}\!\!1&\!\!a^{-1}\\ \!\!0&\!\!1\end{array}\!\!\!\right)$. This element sits in the image of $d_{1}\otimes{\rm id}_{Y_{1}(A^{2})}$. In fact
$(d_{1}\otimes{\rm id}_{Y_{1}(A^{2})})\Big(([g^{-1}_{a}]-[h^{-1}_{a}]-[g^{-1}_{1}]+[h^{-1}_{1}])\otimes(\boldsymbol{\infty},\boldsymbol{0})\Big)=(g^{-1}_{a}-h^{-1}_{a}-g^{-1}_{1}+h^{-1}_{1})[\ ]\otimes(\boldsymbol{\infty},\boldsymbol{0}).$
Finally $({\rm id}_{B_{1}(\mathit{{\rm E}}_{2}(A))}\otimes\partial_{1})\Big(([g^{-1}_{a}]-[h^{-1}_{a}]-[g^{-1}_{1}]+[h^{-1}_{1}])\otimes(\boldsymbol{\infty},\boldsymbol{0})\Big)$ represents $d_{2,0}^{2}(\langle\!\langle a\rangle\!\rangle)$. This is the element
$\bigg(w\Big([g^{-1}_{a}]-[h^{-1}_{a}]-[g^{-1}_{1}]+[h^{-1}_{1}]\Big)-\Big([g^{-1}_{a}]-[h^{-1}_{a}]-[g^{-1}_{1}]+[h^{-1}_{1}]\Big)\bigg)\otimes(\boldsymbol{\infty})\in B_{1}(\mathit{{\rm E}}_{2}(A))\otimes Y_{0}(A^{2}).$
By adding the null element $(d_{2}\otimes{\rm id}_{Y_{0}(A^{2})})(([w|h^{-1}_{a}]-[w|g^{-1}_{a}]+[w|g^{-1}_{1}]-[w|h^{-1}_{1}])\otimes(\boldsymbol{\infty}))$, we see that $d_{2,0}^{2}(\langle\!\langle a\rangle\!\rangle)$ is represented by
$([wg^{-1}_{a}]-[wh^{-1}_{a}]-[wg^{-1}_{1}]+[wh^{-1}_{1}]-[g^{-1}_{a}]+[h^{-1}_{a}]+[g^{-1}_{1}]-[h^{-1}_{1}])\otimes(\boldsymbol{\infty}).$
Now by adding the null element $(d_{2}\otimes{\rm id}_{Y_{0}(A^{2})})(([D(a)^{-1}h^{-1}_{a^{-1}}|wh^{-1}_{a}]-[h^{-1}_{1}|wh^{-1}_{1}])\otimes(\boldsymbol{\infty}))$ (note that $D(z)^{-1}h^{-1}_{z^{-1}}\in\mathrm{B}(A)={\rm Stab}_{\mathit{{\rm E}}_{2}(A)}(\boldsymbol{\infty})$ and $D(z)^{-1}h^{-1}_{a^{-1}}wh^{-1}_{a}=wg^{-1}_{a}$), we see that $d_{2,0}^{2}(\langle\!\langle a\rangle\!\rangle)$ is represented by
$([D(a)^{-1}h^{-1}_{a^{-1}}]-[h^{-1}_{1}]-[g^{-1}_{a}]+[h^{-1}_{a}]+[g^{-1}_{1}]-[h^{-1}_{1}])\otimes(\boldsymbol{\infty}).$
Once more, by adding the null element
$(d_{2}\otimes{\rm id}_{X_{0}(A^{2})})\Big(([h_{1}|-w]-[h_{a^{-1}}|-w])\otimes(\boldsymbol{\infty})\Big)=([g^{-1}_{a}]-[h_{a^{-1}}]-[g^{-1}_{1}]+[h_{1}])\otimes(\boldsymbol{\infty}),$
we see that $d_{2,0}^{2}(\langle\!\langle a\rangle\!\rangle)$ is represented by
$([D(a)^{-1}h^{-1}_{a^{-1}}]-[h^{-1}_{1}]+[h^{-1}_{a}]-[h^{-1}_{1}]-[h_{a^{-1}}]+[h_{1}])\otimes(\boldsymbol{\infty}).$
Note that in $H_{1}(\mathit{{\rm E}}_{2}(A),Y_{0}(A^{2}))$, if $g,h\in\mathrm{B}(A)$, then
$gh]\otimes(\boldsymbol{\infty})=([g]+[h])\otimes(\boldsymbol{\infty})\ \ \text{and}\ \ [h^{-1}]\otimes(\boldsymbol{\infty})=-[h]\otimes(\boldsymbol{\infty}).$
Thus $d_{2,0}^{2}(\langle\!\langle a\rangle\!\rangle)$ in $H_{1}(\mathit{{\rm E}}_{2}(A),Y_{0}(A^{2}))/{\rm im}(d^{1}_{1,1})$ is represented by
$D(a)^{-1}h^{-2}_{a^{-1}}h^{-1}_{a}h^{3}_{1}]\otimes(\boldsymbol{\infty})=\left[\left(\!\begin{array}[]{cc}\!\!a^{-1}&\!\!0\\ \!\!0&\!\!a\end{array}\!\!\!\right)\left(\!\begin{array}[]{cc}\!\!1&\!\!-2a-a^{-1}+3\\ \!\!0&\!\!1\end{array}\!\!\!\right)\right]\otimes(\boldsymbol{\infty}).$
Therefore in $\mathcal{G}_{A}\oplus A_{{A^{\times}}}$ we have
$d_{2,0}^{2}(\langle\!\langle a\rangle\!\rangle)$ $=(\langle a^{-1}\rangle,\overline{-2a-a^{-1}+3})=(\langle a\rangle,\overline{3(1-a)})=3(\langle a\rangle,\overline{1-a}).$
Thus we get the exact sequence
$H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow I(A)/\bar{I}(A)\rightarrow\frac{\mathcal{G}_{A}\oplus A_{A^{\times}}}{\langle(\langle b\rangle,3(b-1))\ |b\in{A^{\times}}\rangle}\rightarrow H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow 0.$
The map
$\frac{A}{M}\rightarrow\frac{\mathcal{G}_{A}\oplus A_{A^{\times}}}{\langle(\langle b\rangle,3(b-1))\ |b\in{A^{\times}}\rangle},\ \ \ \overline{a}\mapsto\overline{(\langle 1\rangle,\bar{a})},$
is an isomorphism with the inverse map
$\frac{\mathcal{G}_{A}\oplus A_{A^{\times}}}{\langle(\langle b\rangle,3(b-1))\ |b\in{A^{\times}}\rangle}\rightarrow\frac{A}{M},\ \ \ \overline{(\langle b\rangle,\bar{a})}\mapsto\overline{a-3b+3}.$
Now (3.2) completes the proof of the theorem. ∎
Corollary 4.2.
If $A$ is universal for ${\rm GE}_{2}$, then $H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\simeq A/M$, where $M$ is the additive subgroup of $A$ generated by $x(a^{2}-1)$ and $3(b+1)(c+1)$ with $x\in A$ and $a,b,c\in{A^{\times}}$.
Proof.
By Theorem 2.5, $H_{1}(Y_{\bullet}(A^{2}))=0$. Now the claim follows from the previous theorem. ∎
Example 4.3.
(i) Let $R$ be a local ring such that its residue field has at least four elements. Then there is $a\in R^{\times}$ such that $a^{2}-1\in R^{\times}$. Now for any $R$-algebra AA and for any $x\in A$, we have
$x=x(a^{2}-1)^{-1}(a^{2}-1)\in M.$
This shows that $A=M$ and therefore by the above theorem
$H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})=0.$
(ii) Let $A$ be a local ring with maximal ideal $\mathfrak{m}_{A}$. Then by the above corollary
$H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\simeq A/M.$
If $|A/\mathfrak{m}_{A}|\geq 4$, then by the above argument we have $H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})=0$. It is not difficult to verify that
[…]
(iii) A commutative semilocal ring $A$ is universal for ${\rm GE}_{2}$ if and only if none of the rings $\mathbb{Z}/2\times\mathbb{Z}/2$ and $\mathbb{Z}/2\times\mathbb{Z}/3$ is a direct factor of $A/J(A)$ [15, Theorem 2.14], where $J(A)$ is the Jacobson radical of $A$. Thus if $A$ is semilocal with the above condition, then by the above corollary
$H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\simeq A/M.$
If $A$ is semilocal such that all its residue fields have more than three elements, i.e. $|A/\mathfrak{m}|>3$ for all $\mathfrak{m}\in{\rm Specm}(A)$, then there is an element $a\in{A^{\times}}$ such that $a^{2}-1\in{A^{\times}}$. Thus $A=M$ and hence
$H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})=0.$
(iv) If $A$ is Artinian, then $A\simeq A_{1}\times\cdots\times A_{r}$, where each $A_{i}$ is an Artinian local ring (see [1, Theorem 8.7]). Hence $\mathit{{\rm E}}_{2}(A)\simeq\mathit{{\rm E}}_{2}(A_{1})\times\cdots\times\mathit{{\rm E}}_{2}(A_{r})$ and thus
$H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\simeq H_{1}(\mathit{{\rm E}}_{2}(A_{1}),\mathbb{Z})\times\cdots\times H_{1}(\mathit{{\rm E}}_{2}(A_{r}),\mathbb{Z}).$
Now using (ii) we can calculate $H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$.
Example 4.4.
(i) For a positive square free integer $d$, let $\mathcal{O}_{-d}$ be the ring of algebraic integers of $\mathbb{Q}(\sqrt{-d})$. Cohn proves that for $d\neq-1,-2,-3,-7,-11$, $\mathcal{O}_{-d}$ is discretely normed [3, §6] and so is universal for ${\rm GE}_{2}$ [3, Theorem 5.2]. Since $\mathcal{O}_{-d}^{\times}=\{1,-1\}$, by the above corollary we have
$H_{1}(\mathit{{\rm E}}_{2}(\mathcal{O}_{-d}),\mathbb{Z})\simeq\mathcal{O}_{-d}/M\simeq\mathbb{Z}\oplus\mathbb{Z}/12.$
On the other hand Swan has shown that
[…]
(see [24, Corollaries 11.2, 12.2, 15.2, 16.2]). Thus the natural map
$H_{1}(\mathit{{\rm E}}_{2}(\mathcal{O}_{-d}),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(\mathcal{O}_{-d}),\mathbb{Z})$
is neither injective nor surjective for $d=5,6$, is injective but not surjective for $d=15$ and is surjective but not injective for $d=19$.
(ii) Let $k$ be a finite field with at least four elements. Then by Example 4.3,
$H_{1}(\mathit{{\rm E}}_{2}(k[X,Y]),\mathbb{Z})=0.$
But it follows from [7, Theorem 1.4] that $H_{1}(\mathit{{\rm SL}}_{2}(k[X,Y]),\mathbb{Z})$ is an infinite group.
(iii) The polynomial ring $A:=\mathbb{Z}[X]$ is universal for ${\rm GE}_{2}$ [3, §2, Theorem 8.2]. Thus by Theorem 2.5 and Lemma 3.1
$H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\simeq H_{1}(\mathit{{\rm E}}_{2}(A),Y_{\bullet}(A^{2}))\simeq H_{1}(\mathit{{\rm SL}}_{2}(A),L_{\bullet}(A^{2})).$
Observe that by the above corollary
$H_{1}(\mathit{{\rm E}}_{2}(\mathbb{Z}[X]),\mathbb{Z})\simeq\mathbb{Z}[X]/M\simeq\mathbb{Z}[X]/12\mathbb{Z}\simeq\mathbb{Z}/12\oplus\bigoplus_{i=1}^{\infty}\mathbb{Z}.$
On the other hand, for any positive integer $m$ there is a surjective homomorphism
$\mathit{{\rm SL}}_{2}(\mathbb{Z}[X])\rightarrow F_{m},$
$F_{m}$ a free group with $m$-generators, such that the group of unipotent matrices $U_{2}(\mathbb{Z}[X])$ is in its kernel. Since for any ring $A$, $\mathit{{\rm E}}_{2}(A)\subseteq U_{2}(A)$, the natural map
$H_{1}(\mathit{{\rm E}}_{2}(\mathbb{Z}[X]),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}[X]),\mathbb{Z})$
has an infinite cokernel.
5. The second homology of $\mathit{{\rm E}}_{2}$
In this section, we study the second homology group of $\mathit{{\rm E}}_{2}(A)$, where $A$ is a commutative ring such that $2\in{A^{\times}}$. The fundamental ideal $I(A)$ plays a central role in this context. More precisely, we investigate a surjective map from $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ onto $I^{2}(A)$.
We have seen in the proof of Theorem 4.1 that the composite $\bar{I}(A)\rightarrow I(A)\overset{d_{2,0}^{2}}{\longrightarrow}\mathcal{G}_{A}\oplus A_{A^{\times}}$, takes $\langle\!\langle a\rangle\!\rangle\in\bar{I}(A)$, $a\in{A^{\times}}$, to the element $(\langle a\rangle,\overline{3(1-a)})$.
Let either $2\in{A^{\times}}$ or $A_{A^{\times}}=0$. In both cases $\overline{3(1-a)}=0$ in $A_{A^{\times}}$. So the above composite sends $\langle\!\langle a\rangle\!\rangle$ to $(\langle a\rangle,0)$. Note that we have a surjective map of abelian groups $A/3A\rightarrow A_{A^{\times}}$. Denote the kernel of $d_{2,0}^{2}:I(A)\rightarrow\mathcal{G}_{A}\oplus A_{A^{\times}}$, by $I^{2}(A)$, i.e.
$I^{2}(A):=\ker(d_{2,0}^{2}).$
We know that $\bar{I}(A)/\bar{I}^{2}(A)\simeq\mathcal{G}_{A}$ [25, Theorem 6.1.11]. Consider the commutative diagram with exact rows
[…]
where the right vertical map is given by $\langle a\rangle\mapsto(\langle a\rangle,0)$. By the Snake Lemma and exact sequence (3.2) we obtain the exact sequence
(5.1) $\bar{I}^{2}(A)\rightarrow I^{2}(A)\rightarrow H_{1}(Y_{\bullet}(A^{2}))_{\mathit{{\rm E}}_{2}(A)}\rightarrow A_{A^{\times}}.$
Lemma 5.1.
If $A_{A^{\times}}=0$, then we have the exact sequence
$\bar{I}^{2}(A)\rightarrow I^{2}(A)\rightarrow H_{1}(Y_{\bullet}(A^{2}))_{\mathit{{\rm E}}_{2}(A)}\rightarrow 0.$
Proof.
This follows immediately from the above exact sequence. ∎
On the other hand, by an easy analysis of the main spectral sequence we have the exact sequence
(5.2) $H_{2}(\mathrm{B}(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A)\rightarrow 0.$
We denote the image of $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in\bar{I}^{2}(A)$ in $I^{2}(A)$ again by $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle$. Now we lift $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in I^{2}(A)$ to an element of $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ through the surjective map
$H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A).$
Proposition 5.2.
Let $A$ be a commutative ring with $2\in{A^{\times}}$. If $a,b\in{A^{\times}}$, then under the surjective map
$H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A),$
the homology class of the cycles $([D(a)|D(b)]+R_{ab}-R_{a}-R_{b}+R_{1})\otimes 1$ maps to $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in I^{2}(A)$, where for any $z\in{A^{\times}}$,
$R_{z}:=$ $[w|g^{-1}_{z}]\!-\![w|h^{-1}_{z}]\!-\![h^{-1}_{z}D(z)^{-1}|wh^{-1}_{z}]\!+\![h_{z^{-1}}|w^{-1}]\!-\![h^{-1}_{z}|D(z)^{-1}]\!+\![D(z)h_{z}|D(z)^{-1}]$
$\!\!\!\!-\![D(z)h_{z}|h^{-1}_{z}]\!+\!2[h^{-1}_{z}|h^{-1}_{z}]\!+\![h^{-2}_{z}|h^{-2}_{z}]\!+\![D(2)h^{-1}_{z}D(2)^{-1}|D(2)]\!-\![D(2)|h^{-1}_{z}]$
with $g_{z}=\left(\!\begin{array}[]{cc}\!\!0&\!\!1\\ \!\!-1&\!\!z\end{array}\!\!\!\right)$ and $h_{z}=\left(\!\begin{array}[]{cc}\!\!1&\!\!z^{-1}\\ \!\!0&\!\!1\end{array}\!\!\!\right)$.
Proof.
We know that $I^{2}(A)\subseteq I(A)\subseteq{\rm GW}(A)=H_{0}(\mathit{{\rm E}}_{2}(A),Z_{1}(A^{2}))$. So we need to study the map
$H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow H_{0}(\mathit{{\rm E}}_{2}(A),Z_{1}(A^{2})).$
From the exact sequence $0\rightarrow Z_{1}^{Y}(A^{2})\rightarrow Y_{1}(A^{2})\rightarrow Y_{0}(A^{2})\rightarrow\mathbb{Z}\rightarrow 0$ we obtain two short exact sequences
$0\rightarrow Z_{1}^{Y}(A^{2})\rightarrow Y_{1}(A^{2})\rightarrow Z_{0}^{Y}(A^{2})\rightarrow 0,\ \ \ \ \ 0\rightarrow Z_{0}^{Y}(A^{2})\rightarrow Y_{0}(A^{2})\rightarrow\mathbb{Z}\rightarrow 0.$
From these we obtain the connecting homomorphisms
$\delta_{1}:H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow H_{1}(\mathit{{\rm E}}_{2}(A),Z_{0}^{Y}(A^{2})),$
$\delta_{2}:H_{1}(\mathit{{\rm E}}_{2}(A),Z_{0}^{Y}(A^{2}))\rightarrow H_{0}(\mathit{{\rm E}}_{2}(A),Z_{1}^{Y}(A^{2})).$
Then
$\delta_{2}\circ\delta_{1}:H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow H_{0}(\mathit{{\rm E}}_{2}(A),Z_{1}^{Y}(A^{2}))$
[…]
$B_{1}(\mathit{{\rm E}}_{2}(A))\!\otimes_{\mathit{{\rm E}}_{2}(A)}\!Z_{0}^{Y}(A^{2})$$B_{1}(\mathit{{\rm E}}_{2}(A))\otimes_{\mathit{{\rm E}}_{2}(A)}\!Y_{1}(A^{2})\!$$B_{0}(\mathit{{\rm E}}_{2}(A))\!\otimes_{\mathit{{\rm E}}_{2}(A)}\!Y_{1}(A^{2})$$B_{0}(\mathit{{\rm E}}_{2}(A))\!\otimes_{\mathit{{\rm E}}_{2}(A)}\!Z_{1}^{Y}(A^{2}).$$\scriptstyle{{\rm id}_{B_{1}}\!\otimes\partial_{1}}$$\scriptstyle{d_{1}\otimes{\rm id}_{Y_{1}}}$$\scriptstyle{{\rm id}_{B_{0}}\otimes{\rm inc}}$
Then
$({\rm id}_{B_{0}}\otimes{\rm inc})\bigg([\ ]\otimes\partial_{2}(X_{z}-X_{1})\bigg)=$ $[\ ]\otimes\bigg((\boldsymbol{0},\boldsymbol{z})-(\boldsymbol{\infty},\boldsymbol{z})-(\boldsymbol{0},\boldsymbol{1})+(\boldsymbol{\infty},\boldsymbol{1})\bigg)$
$=$ $[\ ]\otimes\bigg((g_{z}-h_{z}-g_{1}+h_{1})(\boldsymbol{\infty},\boldsymbol{0})\bigg)$
$=$ $(d_{1}\!\otimes\!{\rm id}_{Y_{1}(A^{2})})\bigg(\!([g^{-1}_{z}]\!-\![h^{-1}_{z}]\!-\![g^{-1}_{1}]\!+\![h^{-1}_{1}])\!\otimes\!(\boldsymbol{\infty},\boldsymbol{0})\!\bigg).$
Thus the element
$U_{z}$ $:=({\rm id}_{B_{1}}\otimes\partial_{1})\bigg(([g^{-1}_{z}]-[h^{-1}_{z}]-[g^{-1}_{1}]+[h^{-1}_{1}])\otimes(\boldsymbol{\infty},\boldsymbol{0})\bigg)$
$=([g^{-1}_{z}]-[h^{-1}_{z}]-[g^{-1}_{1}]+[h^{-1}_{1}])\otimes\partial_{1}(\boldsymbol{\infty},\boldsymbol{0})$
represents an element of $H_{1}(\mathit{{\rm E}}_{2}(A),Z_{0}^{Y}(A^{2}))$, where maps to $\langle\!\langle z\rangle\!\rangle$ through $\delta_{2}$. Now consider the diagram
$B_{2}(\mathit{{\rm E}}_{2}(A))\otimes_{\mathit{{\rm E}}_{2}(A)}\mathbb{Z}$$B_{2}(\mathit{{\rm E}}_{2}(A))\otimes_{\mathit{{\rm E}}_{2}(A)}Y_{0}(A^{2})$$B_{1}(\mathit{{\rm E}}_{2}(A))\otimes_{\mathit{{\rm E}}_{2}(A)}Y_{0}(A^{2})$$B_{1}(\mathit{{\rm E}}_{2}(A))\otimes_{\mathit{{\rm E}}_{2}(A)}Z_{0}^{Y}(A^{2}).$$\scriptstyle{{\rm id}_{B_{2}}\otimes\varepsilon}$$\scriptstyle{d_{2}\otimes{\rm id}_{Y_{0}}}$$\scriptstyle{{\rm id}_{B_{1}}\otimes{\rm inc}}$
Then we have
$({\rm id}_{B_{1}}\otimes{\rm inc})(U_{z})=$ $([g^{-1}_{z}]-[h^{-1}_{z}]-[g^{-1}_{1}]+[h^{-1}_{1}])\otimes\partial_{1}(\boldsymbol{\infty},\boldsymbol{0})$
$=$ $(w-1)([g^{-1}_{z}]-[h^{-1}_{z}]-[g^{-1}_{1}]+[h^{-1}_{1}])\otimes(\boldsymbol{\infty})$
$=$ $-[D(z)]+(d_{2}\otimes{\rm id}_{Y_{0}(A^{2})})\bigg((R_{z}-R_{1})\otimes(\boldsymbol{\infty})\bigg).$
Note that in our calculation we used the formulas
$wg^{-1}_{z}=D(z)^{-1}h^{-1}_{z^{-1}}wh^{-1}_{z},\ \ \ \ \ D(z)h_{u}D(z)^{-1}=h_{u/z^{2}},$
with $u,z\in{A^{\times}}$. Now it is easy to see that
$(d_{2}\!\otimes\!{\rm id}_{Y_{0}(A^{2})})\bigg(([D(a)|D(b)]\!+\!R_{ab}\!-\!R_{a}\!-\!R_{b}\!+\!R_{1})\!\otimes(\boldsymbol{\infty})\bigg)\!=\!({\rm id}_{B_{1}}\otimes{\rm inc})(U_{ab}\!-\!U_{a}\!-\!U_{b}).$
Finally, applying ${\rm id}_{B_{2}}\otimes\varepsilon$ we have
$({\rm id}_{B_{2}}\!\otimes\!\varepsilon)\bigg(\!([D(a)|D(b)]\!+\!R_{ab}\!-\!R_{a}\!-\!R_{b}\!+\!R_{1})\!\otimes\!(\boldsymbol{\infty})\!\bigg)\!=\!([D(a)|D(b)]\!+\!R_{ab}\!-\!R_{a}\!-\!R_{b}\!+\!R_{1})\!\otimes\!1.$
The element $([D(a)|D(b)]+R_{ab}-R_{a}-R_{b}+R_{1})\otimes 1$ is a cycle and represents an element of $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ which maps to $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in I^{2}(A)$. This completes the proof of the proposition. ∎
Remark 5.3.
In [10, §7] Hutchinson has constructed a similar cycles as in the above proposition over fields. Let $F$ be a field with at least four elements. Let $a,b\in F^{\times}$ and $\lambda\in F^{\times}$ such that $\lambda^{2}-1\in F^{\times}$. Then Hutchinson has constructed a cycle
$F(a,b)_{\lambda}\in B_{2}(\mathit{{\rm SL}}_{2}(F))\otimes_{\mathit{{\rm SL}}_{2}(F)}\mathbb{Z}$
such that its homology class in $H_{2}(\mathit{{\rm SL}}_{2}(F),\mathbb{Z})$ is independent of the choice of $\lambda$ and maps to $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in I^{2}(F)$ (see [10, Proposition 7.1, Lemma 7.5, Theorem 7.8]). (Note that the existence of $\lambda\in{A^{\times}}$ such that $\lambda^{2}-1\in A^{\times}$ implies that $A_{A^{\times}}=0$.)
Remark 5.4.
Let $A$ be a semilocal ring such that $2\in{A^{\times}}$ and all its residue fields have more that four elements. Then the group $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ is generated by the homology class of the cycles $([D(a)|D(b)]+R_{ab}-R_{a}-R_{b}+R_{1})$, ([D(a)|D(b)][D(b)|D(a)])([D(a)|D(b)]-[D(b)|D(a)]), $a,b\in{A^{\times}}$, and $([E_{12}(x)|E_{12}(y)]-[E_{12}(y)|E_{12}(x)])$, $x,y\in A$:
In fact since $A$ is universal for ${\rm GE}_{2}$ (see Example 4.3), by Proposition 2.5, $H_{1}(Y_{\bullet}(A^{2}))=0$. Thus by Lemma 5.1, the map $\bar{I}^{2}(A)\rightarrow I^{2}(A)$ is surjective. This shows that $I^{2}(A)$ is generated by the elements of the form $\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle$, $a,b\in{A^{\times}}$. Now consider the exact sequence
$H_{2}(\mathrm{B}(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\overset{\theta}{\rightarrow}I^{2}(A)\rightarrow 0.$
Let $X\in H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ and let $\theta(X)=\sum\langle\!\langle a_{i}\rangle\!\rangle\langle\!\langle b_{i}\rangle\!\rangle$. If $F(a,b):=([D(a)|D(b)]+R_{ab}-R_{a}-R_{b}+R_{1})$, then $\theta(X-\sum\overline{F(a_{i},b_{i})})=0$. It follows from this that $X-\sum\overline{F(a_{i},b_{i})}$ sits in the image of $H_{2}(\mathrm{B}(A),\mathbb{Z})$. Consider the extension $1\rightarrow\mathrm{N}(A)\rightarrow\mathrm{B}(A)\rightarrow\mathrm{T}(A)\rightarrow 1$, where $\mathrm{N}(A):=\bigg\{\begin{pmatrix}1&b\\ 0&1\end{pmatrix}:a\in{A^{\times}},b\in A\bigg\}$. Since any residue field of $A$ has more than four elements, there is $\lambda\in{A^{\times}}$ such that $\lambda^{2}-1\in{A^{\times}}$. This implies that $A_{A^{\times}}=0$. It follows from this that for any $k\geq 0$, $H_{k}({A^{\times}},A)=0$ [17, Corollary 3.2]. By studying the Lyndon/Hochschild-Serre spectral sequence of this extension we obtain the exact sequence
$H_{2}(A,\mathbb{Z})_{A^{\times}}\rightarrow H_{2}(\mathrm{B}(A),\mathbb{Z})\rightarrow H_{2}(\mathrm{T}(A),\mathbb{Z})\rightarrow 0.$
Note that we have the splitting map ${\rm inc}_{\ast}:H_{2}(\mathrm{T}(A),\mathbb{Z})\rightarrow H_{2}(\mathrm{B}(A),\mathbb{Z})$. The image of $H_{2}(A,\mathbb{Z})\simeq A\wedge A$ in $H_{2}(\mathrm{B}(A),\mathbb{Z})$ is generated by the homology classes of the cycles
$G(x,y):=([E_{12}(x)|E_{12}(y)]-[E_{12}(y)|E_{12}(x)])\otimes 1,\ \ \ x,y\in A.$
Moreover, $H_{2}(\mathrm{T}(A),\mathbb{Z})\simeq\mathrm{T}(A)\wedge\mathrm{T}(A)$ is generated by the homology classes of the cycles
$H(a,b):=([D(a)|D(b)]-[D(b)|D(a)])\otimes 1.$
Now the claim follows from the fact that $\ker(\theta)$ is generated by the images of the homology classes of the cycles $G(x,y)$ and $H(a,b)$. This proves our claim.
6. The third homology of $\mathit{{\rm E}}_{2}$
A potential refined Bloch–Wigner exact sequence is an exact sequence that describes the structure of the third homology group of $\mathit{{\rm E}}_{2}(A)$. Such an exact sequence has been established in a few cases [8], [11], [18]. We believe that in the present section we take a significant step toward a better understanding of such an exact sequence over semilocal rings.
Let $\overline{\mathcal{RP}}(A)$ be the quotient of the free $\mathbb{Z}[\mathcal{G}_{A}]$-module generated by symbols $[a]$, $a\in\mathcal{W}_{A}$, by the $\mathbb{Z}[\mathcal{G}_{A}]$-submodule generated by the elements
$a]-[b]+\langle a\rangle\bigg[\frac{b}{a}\bigg]-\langle a^{-1}-1\rangle\Bigg[\frac{1-a^{-1}}{1-b^{-1}}\Bigg]+\langle 1-a\rangle\Bigg[\frac{1-a}{1-b}\Bigg],$
where $a,b,a/b\in\mathcal{W}_{A}$. By a direct computation one can show that the map
$\bar{\lambda}_{1}:\overline{\mathcal{RP}}(A)\rightarrow\mathcal{I}_{A}^{2}\subseteq\mathbb{Z}[\mathcal{G}_{A}],\ \ \ \ \ \ [a]\mapsto\langle\!\langle a\rangle\!\rangle\langle\!\langle 1-a\rangle\!\rangle,$
is a well-defined $\mathbb{Z}[\mathcal{G}_{A}]$-homomorphism. Let
$S_{\mathbb{Z}}^{2}({A^{\times}}):=({A^{\times}}\otimes_{\mathbb{Z}}{A^{\times}})/\langle a\otimes b+b\otimes a:a,b\in{A^{\times}}\rangle.$
If we consider $S_{\mathbb{Z}}^{2}({A^{\times}})$ as a trivial $\mathcal{G}_{A}$-module, then
$\bar{\lambda}_{2}:\overline{\mathcal{RP}}(A)\rightarrow S_{\mathbb{Z}}^{2}({A^{\times}}),\ \ \ \ \ [a]\mapsto a\otimes(1-a),$
is a homomorphism of $\mathbb{Z}[\mathcal{G}_{A}]$-modules. Hutchinson defined the $\mathbb{Z}[\mathcal{G}_{A}]$-modules $\overline{\mathcal{RP}}_{1}(A)$ and $\overline{\mathcal{RB}}(A)$ as follows
$\overline{\mathcal{RP}}_{1}(A):=\ker(\bar{\lambda}_{1}:\overline{\mathcal{RP}}(A)\rightarrow\mathcal{I}_{A}^{2}),$
$\overline{\mathcal{RB}}(A):=\ker(\bar{\lambda}_{2}|_{\overline{\mathcal{RP}}_{1}(A)}:\mathcal{RP}_{1}(A)\rightarrow S_{\mathbb{Z}}^{2}({A^{\times}}))$
(see [8, page 28], [9, Subsection 2.3]).
From the complex $Y_{4}(A^{2})\rightarrow Y_{3}(A^{2})\rightarrow Z_{2}^{Y}(A^{2})\rightarrow 0$ we obtain the complex of $\mathcal{G}_{A}$-modules
$Y_{4}(A^{2})_{\mathit{{\rm E}}_{2}(A)}\rightarrow Y_{3}(A^{2})_{\mathit{{\rm E}}_{2}(A)}\rightarrow Z_{2}^{Y}(A^{2})_{\mathit{{\rm E}}_{2}(A)}\rightarrow 0.$
We have seen that $Y_{3}(A^{2})_{\mathit{{\rm E}}_{2}(A)}$ is a free $\mathbb{Z}[\mathcal{G}_{A}]$-module generated by the symbols $[x]$, $x\in\mathcal{W}_{A}$. The orbits of the action of $\mathit{{\rm E}}_{2}(A)$ $Y_{4}(A)$ are represented by
$\langle a\rangle[x,y]:=({\boldsymbol{\infty}},{\boldsymbol{0}},{\boldsymbol{a}},\boldsymbol{ax},\boldsymbol{ay}),\ \ \langle a\rangle\in\mathcal{G}_{A},x,y,x/y\in\mathcal{W}_{A}.$
Thus $Y_{4}(A^{2})_{\mathit{{\rm E}}_{2}(A)}$ is the free $\mathbb{Z}[\mathcal{G}_{A}]$-module generated by the symbols $[x,y]$, $x,y,x/y\in\mathcal{W}_{A}$. It is straightforward to check that
$\overline{\partial_{4}}([x,y])=[x]-[y]+\langle x\rangle\bigg[\frac{y}{x}\bigg]-\langle x^{-1}-1\rangle\Bigg[\frac{1-x^{-1}}{1-y^{-1}}\Bigg]+\langle 1-x\rangle\Bigg[\frac{1-x}{1-y}\Bigg].$
Thus we obtain a natural map
$\eta:\overline{\mathcal{RP}}(A)\rightarrow H_{0}(\mathit{{\rm E}}_{2}(A),Z_{2}^{Y}(A^{2})).$
If $Y_{\bullet}(A)\rightarrow\mathbb{Z}$ is exact in dimension <4<4, then the above map becomes an isomorphism.
Following Coronado and Hutchinson [5, §3] we define the refined scissors congruence group of $A$ as follows:
$\mathcal{RP}(A):=H_{0}(\mathit{{\rm SL}}_{2}(A),Z_{2}^{L}(A^{2})).$
It follows from Lemma 2.6, that
$\mathcal{RP}(A)\simeq H_{0}(\mathit{{\rm E}}_{2}(A),Z_{2}^{Y}(A^{2}))=Z_{2}^{Y}(A^{2})_{\mathit{{\rm E}}_{2}(A)}.$
The inclusion $Z_{2}^{Y}(A^{2})\rightarrow Y_{2}(A^{2})$ induces the natural map
$\lambda_{1}:={\rm inc}:\mathcal{RP}(A)=Z_{2}^{Y}(A^{2})_{\mathit{{\rm E}}_{2}(A)}\rightarrow Y_{2}(A^{2})_{\mathit{{\rm E}}_{2}(A)}\simeq\mathbb{Z}[\mathcal{G}_{A}].$
It is straightforward to check that $\bar{\lambda}_{1}=\lambda_{1}\circ\eta$. We denote the kernel of $\lambda_{1}$ by $\mathcal{RP}_{1}(A)$ and call it the refined scissors congruence group of $A$. Note that we have a natural map
$\overline{\mathcal{RP}}_{1}(A)\rightarrow\mathcal{RP}_{1}(A).$
Let $Y_{\bullet}(A)$ be exact in dimension $1$. From the short exact sequence
$0\rightarrow Z_{2}^{Y}(A^{2})\rightarrow Y_{2}(A^{2})\overset{\partial_{2}^{Y}}{\rightarrow}Z_{1}^{Y}(A^{2})\rightarrow 0,$
we obtain the (long) exact sequence
$H_{1}(\mathit{{\rm E}}_{2}(A),Y_{2}(A^{2}))\overset{\bar{\partial}_{2}^{Y}}{\rightarrow}H_{1}(\mathit{{\rm E}}_{2}(A),Z_{1}^{Y}(A^{2}))\rightarrow\mathcal{RP}(A)\overset{\lambda_{1}}{\longrightarrow}\mathbb{Z}[\mathcal{G}_{A}].$
Since
$H_{1}(\mathit{{\rm E}}_{2}(A),Y_{2}(A^{2}))\simeq H_{1}\bigg(\mathit{{\rm E}}_{2}(A),\bigoplus_{\langle a\rangle\in\mathcal{G}_{A}}{\rm Ind}_{\mu_{2}(A)}^{\mathit{{\rm E}}_{2}(A)}\mathbb{Z}\bigg)\simeq\mu_{2}(A)\otimes_{\mathbb{Z}}\mathbb{Z}[\mathcal{G}_{A}],$
we have the exact sequence
$\mu_{2}(A)\otimes_{\mathbb{Z}}\mathbb{Z}[\mathcal{G}_{A}]\overset{\bar{\partial}_{2}}{\rightarrow}\mathrm{E}_{2,1}^{1}\rightarrow\mathcal{RP}_{1}(A)\rightarrow 0.$
From the commutative diagram
$Y_{2}(A^{2})$$Z_{1}^{Y}(A^{2})$$Y_{1}(A^{2})$$Y_{1}(A^{2})$$\scriptstyle{\partial_{2}^{Y}}$$\scriptstyle{\partial_{2}}$$\scriptstyle{{\rm inc}}$
we obtain the commutative diagram
$\mu_{2}(A)\otimes_{\mathbb{Z}}\mathbb{Z}[\mathcal{G}_{A}]$$\mathrm{E}_{2,1}^{1}$$\mathcal{RP}_{1}(A)\longrightarrow 0$$\mu_{2}(A)$$\mu_{2}(A).$$\scriptstyle{\bar{\partial}_{2}}$$\scriptstyle{d_{2,1}^{1}}$
It is not difficult to see that $(d_{2,1}^{1}\circ\bar{\partial}_{2}^{Y})(b\otimes\langle a\rangle)=b$. Thus from the above commutative diagram we obtain the exact sequence
$\mu_{2}(A)\otimes_{\mathbb{Z}}\mathcal{I}_{A}\rightarrow\mathrm{E}_{2,1}^{2}\rightarrow\mathcal{RP}_{1}(A)\rightarrow 0.$
Consider the differential $d_{2,1}^{2}:E_{2,1}^{2}\rightarrow H_{2}(\mathrm{B}(A),\mathbb{Z})$. In the rest of this section we will assume that the surjective map $p:\mathrm{B}(A)\rightarrow\mathrm{T}(A)$ induces the isomorphism
$H_{2}(\mathrm{B}(A),\mathbb{Z})\overset{p_{\ast}}{\simeq}H_{2}(\mathrm{T}(A),\mathbb{Z}).$
For example over semilocal rings we have the following result.
Proposition 6.1.
Let $A$ be a semilocal ring. For any maximal ideal $\mathfrak{m}\in{\rm Specm}(A)$, let either $A/\mathfrak{m}$ be infinite or if $|A/\mathfrak{m}|=p^{d}$, then (p1)d>2(n+1)(p-1)d>2(n+1) ((p1)d>2n(p-1)d>2n when $A$ is a domain). Then $H_{k}({\rm T}(A),\mathbb{Z})\simeq H_{k}({\rm B}(A),\mathbb{Z})$ for all $k\leq n$.
Proof.
If $A$ is a domain, then the proof is very similar to the proof of a similar result over local domains presented in [11, §3]. Just there one should replace the maximal ideal of the local domain (denoted by $\mathcal{M}$ in [11, §3]) with the Jacobson radical $J(A)$ of $A$.
Now let $A$ be a general semilocal ring with the given condition. Then similar to the proof of [11, Proposition 3.19], we can show that $H_{n}(\mathrm{T}(A),k)\simeq H_{n}(\mathrm{B}(A),k)$, where $k$ is a prime field and (p1)d>2n(p-1)d>2n. Now the claim follows from [16, Lemma 2.3]. ∎
Thus let $H_{2}(\mathrm{B}(A),\mathbb{Z})\overset{p_{\ast}}{\simeq}H_{2}(\mathrm{T}(A),\mathbb{Z})$. Then the composite
$\mu_{2}(A)\otimes_{\mathbb{Z}}\mathcal{I}_{A}\rightarrow\mathrm{E}_{2,1}^{2}\overset{d_{2,1}^{2}}{\longrightarrow}H_{2}(\mathrm{B}(A),\mathbb{Z})\simeq H_{2}(\mathrm{T}(A),\mathbb{Z})\simeq{A^{\times}}\wedge{A^{\times}}$
is given by $b\otimes\langle\!\langle a\rangle\!\rangle\mapsto b\wedge a$ (see [17, Lemma 4.1]). Thus we have the commutative diagram
[…]
Consider the injective map
$\alpha:\frac{{A^{\times}}\wedge{A^{\times}}}{{A^{\times}}\wedge\mu_{2}(A)}\rightarrow S_{\mathbb{Z}}^{2}({A^{\times}}),\ \ \ \ a\wedge b\mapsto 2(a\otimes b).$
One can show that the composite map
$\overline{\mathcal{RP}}_{1}(A)\rightarrow\mathcal{RP}_{1}(A)\rightarrow\frac{{A^{\times}}\wedge{A^{\times}}}{\mu_{2}(A)\wedge{A^{\times}}}\rightarrow S_{\mathbb{Z}}^{2}({A^{\times}})$
coincides with $\bar{\lambda}_{2}|_{\overline{\mathcal{RP}}_{1}(A)}:\overline{\mathcal{RP}}_{1}(A)\rightarrow S_{\mathbb{Z}}^{2}({A^{\times}})$. We denote the kernel of
$\mathcal{RP}_{1}(A)\rightarrow\frac{{A^{\times}}\wedge{A^{\times}}}{\mu_{2}(A)\wedge{A^{\times}}}\rightarrow S_{\mathbb{Z}}^{2}({A^{\times}})$
with $\mathcal{RB}(A)$ and call it the refined Bloch group of $A$. Note that we have a natural map
$\overline{\mathcal{RB}}(A)\rightarrow\mathcal{RB}(A).$
Now from the above commutative diagram we obtain a surjective map $\mathrm{E}_{2,1}^{\infty}\twoheadrightarrow\mathcal{RB}(A)$. By an easy analysis of the main spectral sequence we obtain the surjective map
$H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow\mathrm{E}_{2,1}^{\infty}.$
Combining these two maps we obtain a surjective map
$H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow\mathcal{RB}(A).$
This map factors through $H_{3}(\mathit{{\rm PE}}_{2}(A),\mathbb{Z})$, where
$\mathit{{\rm PE}}_{2}(A):=\mathit{{\rm E}}_{2}(A)/\{\pm I_{2}\}.$
We have proved the following version of a refined Bloch-Wigner exact sequence in [19].
Theorem 6.2.
Let $A$ be a semilocal ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu(A)\rightarrow\mu(F)$ is injective. For any maximal ideal $\mathfrak{m}\in{\rm Specm}(A)$ let either $A/\mathfrak{m}$ be infinite or if $|A/\mathfrak{m}|=p^{d}$, then (p1)d>8(p-1)d>8 ((p1)d>6(p-1)d>6 when $A$ is a domain). If either $1\in{A^{\times}}^{2}$ or $|\mathcal{G}_{A}|\leq 4$, then we have the exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(A),\widetilde{\mu}(A))\rightarrow H_{3}(\mathit{{\rm PE}}_{2}(A),\mathbb{Z})\rightarrow\mathcal{RB}(A)\rightarrow 0,$
where $\widetilde{\mu}(A)=\mu(A)/\mu_{2}(A)$.
Proof.
See [19, Theorem 3.1, Theorem 3.2, Corollary 3.3] and use Proposition 6.1 and [19, Lemma 2.8]. ∎
Let $\mathcal{A}$ be a finite cyclic group. If $2\mid|\mathcal{A}|$, let $\mathcal{A}^{\sim}$ denote the unique non-trivial extension of $\mathcal{A}$ by $\mathbb{Z}/2$. If $2\nmid|\mathcal{A}|$, we define $\mathcal{A}^{\sim}:=\mathcal{A}$. Thus if $n=|\mathcal{A}|$, then
$\mathcal{A}^{\sim}\simeq\mathbb{Z}/\gcd(n^{2},2n).$
Let $\mathcal{A}$ be an ind-cyclic group, i.e. $\mathcal{A}$ is direct limit of its finite cyclic subgroups. Then by passing to the limit we can define $\mathcal{A}^{\sim}$. Note that if $\mathcal{A}$ has no $2$-torsion, then $\mathcal{A}^{\sim}=\mathcal{A}$. There is always a canonical injective homomorphism $\mathcal{A}\hookrightarrow\mathcal{A}^{\sim}$ whose composition with the projection $\mathcal{A}^{\sim}\rightarrow\mathcal{A}$ coincides with multiplication by $2$.
The connection between the third homology groups of $\mathit{{\rm E}}_{2}(A)$ and $\mathit{{\rm PE}}_{2}(A)$ is given in [20, Proposition 5.1].
Theorem 6.3.
Let $A$ be a semilocal ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu_{2}(A)\simeq\mu_{2}(F)$. If all the residue fields of $A$ have more than three elements, then we have the exact sequence
$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}\rightarrow H_{3}(\mathit{{\rm PE}}_{2}(A),\mathbb{Z})\rightarrow 0,$
where the map $\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ is induced by the product map $\mu_{2}(A)\times\mathit{{\rm E}}_{2}(A)\rightarrow\mathit{{\rm E}}_{2}(A)$, $(-1,X)\mapsto-X$.
Proof.
We may assume that ${\rm char}(F)\neq 2$. Since for any maximal ideal $\mathfrak{m}\in{\rm Specm}(A)$, $|A/\mathfrak{m}|\geq 4$, there is $a\in{A^{\times}}$ such that $a^{2}-1\in{A^{\times}}$. It follows from this that $M=A$. Thus by Theorem 4.1 we have $H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})=0$. Now by [20, Theorem 3.1] we obtain the desired exact sequence. ∎
The above two theorems are results of our attempt to find an answer to the following question of Coronado and Hutchinson [5, page 3].
Question 6.4.
Is it true that for any local domain $A$ one has the exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow\mathcal{RB}(A)\rightarrow 0?$
The main result of this section is as follows, which generalizes the main result of [18, Theorem 5.1, Corollary 5.2].
Theorem 6.5.
Let $A$ be a semilocal ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu(A)\rightarrow\mu(F)$ is injective. For any maximal ideal $\mathfrak{m}\in{\rm Specm}(A)$ let either $A/\mathfrak{m}$ be infinite or if $|A/\mathfrak{m}|=p^{d}$, then (p1)d>8(p-1)d>8 ((p1)d>6(p-1)d>6 when $A$ is a domain). If either $1\in{A^{\times}}^{2}$ or $|\mathcal{G}_{A}|\leq 4$, then we have the exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)}\rightarrow\mathcal{RB}(A)\rightarrow 0,$
where $(-1)\otimes\langle\!\langle a\rangle\!\rangle\langle\!\langle b\rangle\!\rangle\in\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)$ maps in $H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ to the class of the cycle which is obtained from the shuffle product of $1]\in B_{1}(\mu_{2}(A))_{\mu_{2}(A)}$ with the cycle $D(a)|D(b)]+R_{ab}-R_{a}-R_{b}+R_{1}\in B_{2}(\mathit{{\rm E}}_{2}(A))_{\mathit{{\rm E}}_{2}(A)}$.
Proof.
Consider the surjective maps
$\frac{H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}\twoheadrightarrow H_{3}(\mathit{{\rm PE}}_{2}(A),\mathbb{Z})\twoheadrightarrow\mathcal{RB}(A).$
Let $\mathcal{K}$ be the kernel of this composite. Then we have the commutative diagram with exact rows
[…]
(see Theorem 6.2). From this diagram and Theorem 6.3 we obtain the exact sequence
$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\mathcal{K}\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(A),\widetilde{\mu}(A))\rightarrow 0.$
This clearly implies that $\mathcal{K}\simeq{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}$ and thus we have the exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}\rightarrow\mathcal{RB}(A)\rightarrow 0.$
By Proposition 6.1, $H_{2}({\rm B}(A),\mathbb{Z})\simeq H_{2}({\rm T}(A),\mathbb{Z})$. Under the composition
${\rm T}(A)\wedge{\rm T}(A)\overset{\simeq}{\longrightarrow}H_{2}({\rm T}(A),\mathbb{Z})\overset{\simeq}{\longrightarrow}H_{2}({\rm B}(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$
$D(a)\wedge D(b)$ maps to the homology class of the cycle ([D(a)|D(b)][D(b)|D(a)])([D(a)|D(b)]-[D(b)|D(a)]), $a,b\in{A^{\times}}$. Now from the commutative diagram
$\mu_{2}(A)\otimes_{\mathbb{Z}}({\rm T}(A)\wedge{\rm T}(A))$$\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}({\rm T}(A),\mathbb{Z})$$\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$${\rm T}(A)\wedge{\rm T}(A))\wedge{\rm T}(A)$$H_{3}({\rm T}(A),\mathbb{Z})$$H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$}
it follows that the right vertical map is given by
$(-1)\otimes(D(a)\wedge D(b))\mapsto D(-1)\wedge D(a)\wedge D(b).$
Note that here we denoted the image of $D(-1)\wedge D(a)\wedge D(b)\in\bigwedge_{\mathbb{Z}}^{3}\mathrm{T}(A)$ in $H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ again by $D(-1)\wedge D(a)\wedge D(b)$. If $\mathfrak{i}^{2}=-1$, then
$D(-1)\wedge D(a)\wedge D(b)=2(D(\mathfrak{i})\wedge D(a)\wedge D(b))=0.$
Let $|\mathcal{G}_{A}|\leq 4$. If $1\in{A^{\times}}^{2}$, we proceed as above. If $1\notin{A^{\times}}^{2}$, then $\mathcal{G}_{A}=\{1,-1,u,-u\}$ for some $u\in{A^{\times}}$. So we may take $a,b\in\{1,-1,u,-u\}$ and in any case
$D(-1)\wedge D(a)\wedge D(b)=0.$
Thus $\frac{H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}=\frac{H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)}$. This completes the proof of the theorem. ∎
In order to have a positive answer to Question 6.4, we should have a positive answer to the following question, which was raised in [20].
Question 6.6.
Let $A$ be a local domain such that $|A/\mathfrak{m}_{A}|\neq 2$. Is the product map
$\rho_{\ast}:\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$
trivial?
We do not know the answer of the above question even over a general infinite field. For more on this we refer the interested readers to [20, Example 4.4, Remark 5.2, Proposition 5.4].
Remark 6.7.
Let $A$ be a local domain such that $|A/\mathfrak{m}_{A}|\neq 2,3,4,5,7,8,9,16,27,32,64$. Then the results of this paper suggest that to give a positive answer to Question 6.4, one may show that
(i) we have the projective Bloch-Wigner exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(A),\widetilde{\mu}(A))\rightarrow H_{3}(\mathit{{\rm PE}}_{2}(A),\mathbb{Z})\rightarrow\mathcal{RB}(A)\rightarrow 0,$
(ii) the product map $\rho_{\ast}:\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ is trivial.
Remark 6.8.
Although significant progress has been made toward understanding the low-dimensional homology of $\mathit{{\rm E}}_{2}(A)$ (see the list of references below), many problems remain open for future research.
For the first homology, we hope to obtain a deeper understanding of the kernel of the surjective map $A/M\rightarrow H_{1}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ (from Theorem 4.1) for an arbitrary commutative ring. This requires a more detailed analysis of the structure of the group $H_{1}(Y_{\bullet})$.
Although we have a reasonable understanding of the second homology of $\mathit{{\rm E}}_{2}(A)$ over semilocal rings, determining the precise structure of this group over general rings remains an open problem. For results concerning the structure of the group $H_{2}(\mathit{{\rm E}}_{2}(A),\mathbb{Z})$ for $A=\mathbb{Z}[1/n]$, see [10] and [21] and for $A$ a discrete valuation rings see [13].
It appears that the third homology of $\mathit{{\rm E}}_{2}(A)$ over semilocal rings is best described by a refined Bloch–Wigner exact sequence. As explained in the previous remark, a positive answer to Problem 6.4 depends on a positive answer to Problem 6.6 and on the existence of a projective refined Bloch–Wigner exact sequence. We believe that both questions should have affirmative answers, provided that the residue fields of the local rings are sufficiently large. It is worth noting that even over a general infinite field, a definitive result is still lacking.